OpenStudy (narissa):

help 1/4x-8+2/4x=19 x=

OpenStudy (narissa):

@Rushwr

OpenStudy (narissa):

@texaschic101

OpenStudy (narissa):

@mathstudent55

OpenStudy (narissa):

@IrishBoy123

OpenStudy (narissa):

@ChantySquirrel1129**

OpenStudy (rushwr):

This equation is a little tricky. The way it is written here is confusing.... Can u rewrite it using "equation"

OpenStudy (chantysquirrel1129**):

\[\frac{ 1 }{ 4 }x - 8 + \frac{ 2 }{ 4 }x=19\] I think she meant that, I'm not really sure :p c:

OpenStudy (narissa):

@Rushwr

OpenStudy (rushwr):

\[\frac{ 1 }{ 4}x -8 +\frac{ 2 }{ 4 }x =19\] \[\frac{ 1 }{ 4}x + \frac{ 2 }{ 4 }x = 19+8 = 27\] \[\frac{ 3 }{ 4 }x = 27\] \[x = \frac{ 4*27 }{ 3 } = 36\]

OpenStudy (narissa):

thank u sooo much

OpenStudy (narissa):

i have 3 more can u help